### v1

# An Empirical Study of Federated Prompt Learning for Vision Language Model

Abstract
The Vision Language Model (VLM) excels in aligning vision and language representations, and prompt learning has emerged as a key technique for adapting such models to downstream tasks. However, the application of prompt learning with VLM in federated learning (FL) scenarios remains underexplored. This paper systematically investigates the behavioral differences between language prompt learning (LPT) and vision prompt learning (VPT) under data heterogeneity challenges, including label skew and domain shift. We conduct extensive experiments to evaluate the impact of various FL and prompt configurations, such as client scale, aggregation strategies, and prompt length, to assess the robustness of Federated Prompt Learning (FPL). Furthermore, we explore strategies for enhancing prompt learning in complex scenarios where label skew and domain shift coexist, including leveraging both prompt types when computational resources allow. Our findings offer practical insights into optimizing prompt learning in federated settings, contributing to the broader deployment of VLMs in privacy-preserving environments.
1 Introduction
Figure 1: Background for Federated Prompt Learning. Participants learn shared visual or textual prompts. However, under data heterogeneity (Label Skew and Domain Shift), aggregated prompts struggle to capture consistency among clients. See details in Sec. 1.
The advent of large-scale Vision Language Models, such as Contrastive Language-Image Pretraining (CLIP) Radford et al. (2021), has redefined multi-modal learning by enabling effective alignment between visual and textual representations. These models have demonstrated remarkable adaptability across various downstream tasks with prompt tuning Shin et al. (2020); Lester et al. (2021), particularly in zero-shot and few-shot settings. However, due to the inconsistency of upstream and downstream data distribution and training costs, large-scale Vision Language Models such as CLIP still face many challenges in training and deployment.
Prompt learning Bang et al. (2024); Li et al. (2024b), a lightweight yet effective paradigm for fine-tuning large pre-trained models, has emerged as a promising solution. By introducing task-specific prompts rather than retraining full model parameters, prompt learning offers an efficient means of adapting models to new tasks. While this approach has been extensively examined in centralized settings, its potential for optimizing VLM under federated conditions remains underexplored Guo et al. (2023b). In these scenarios, the design and tuning of prompts are particularly critical, as they must account for the decentralized nature of data and the limited communication resources available in FL.
In Federated Learning scenarios, the challenges of data heterogeneity Zhao et al. (2018); Kairouz et al. (2019); Li et al. (2020a); Tan et al. (2023) are further compounded by issues like Label Skew and Domain Shift as shown in Fig. 1. Label skew occurs when different clients have significantly imbalanced class distributions, leading to poor generalization on unseen data. Domain shift arises when the underlying data distributions vary across clients, making it difficult for the model to converge to a globally optimal solution. For VLMs enhanced with prompt learning, these issues manifest in distinct ways: language prompts may be less effective when the semantic representation of classes varies widely due to label skew, whereas visual prompts may struggle to capture consistent image features across diverse domains.
Motivated by these observations, our work conducts an in-depth empirical study of prompt learning for VLM in FL settings. We systematically analyze the distinct behaviors of language prompt learning and vision prompt learning in scenarios characterized by Label Skew and Domain Shift. Specifically, we explore how each type of prompt responds to imbalanced class distributions and varying domain characteristics, thereby uncovering the strengths and limitations inherent in each approach. We study the influence of different FL configurations alongside various prompt settings. This analysis aims to explore the robustness of federated prompt learning. Furthermore, recognizing that real-world applications often involve simultaneous challenges from both Label Skew and Domain Shift, we investigate strategies to boost the effectiveness of prompt learning in such complex environments.
Our investigation is structured around three main research objectives:
  • [Uncaptioned image]
    Q1: Do Language Prompt Tuning (LPT) and Vision Prompt Tuning (VPT) show Behavior Discrepancies under label skew and domain shift?
  • [Uncaptioned image]
    Q2: Does federated prompt learning exhibits Robustness Patterns in various federated settings and prompt configurations?
  • [Uncaptioned image]
    Q3: Do federated visual and textual prompts appear Collaboration Effect for complex heterogeneity?
By addressing these three key areas, our work not only elucidates the differential impacts of language and vision prompts in federated settings but also provides practical guidelines for optimizing prompt learning in large-scale, distributed vision-language models. This study paves the way for more robust and resource-efficient deployments of VLM in real-world FL applications.
2 Related Work
2.1 Federated Learning
Federated Learning (FL) enables collaborative model training across decentralized clients while preserving data privacy Konečný et al. (2016); Kairouz et al. (2019); Li et al. (2020b); Huang et al. (2022); Dai et al. (2023); Huang et al. (2023a, 2024). However, it faces significant challenges, particularly in scenarios involving label skew and domain shift, which substantially degrade model performance due to the presence of non-IID (non-independent and identically distributed) data Luo et al. (2021); Li et al. (2022); Ma et al. (2022); Ye et al. (2023); Dai et al. (2023); Hu et al. (2024); Wang et al. (2025). Label skew occurs when the class distributions vary significantly across clients, leading to poor generalization on unseen data. Domain shift, on the other hand, arises when the input data distribution differs across clients, further complicating model convergence.
To address label skew, FedProx Li et al. (2020c)introduced a proximal term to the optimization objective, stabilizing training in non-IID settings. Similarly, MOON Li et al. (2021a) proposed contrastive learning at the client level to align model representations and mitigate discrepancies caused by skewed labels. For domain shift, FedBN Li et al. (2021b) employed client-specific batch normalization layers, allowing local domain-specific feature extraction without compromising global model updates. Other methods, such as MFLLiu et al. (2020) and CFLSattler et al. (2021), attempted to balance global and local objectives by clustering clients with similar data distributions or introducing momentum-based updates to improve convergence under domain shifts. These methods focus on improving the communication efficiency and convergence of FL systems when clients experience heterogeneous data distributions. Despite these advancements in handling label skew and domain shift individually, there is a need for further research on how large-scale models like CLIP can be better adapted and applied in FL settings when facing these challenges. This paper aims to explore how CLIP can leverage prompt learning to maintain robust performance in federated settings, ensuring better cross-client adaptation and efficient model convergence despite the inherent challenges of data distribution.
2.2 Prompt Learning
Prompt learning is a powerful and flexible approach for adapting pre-trained models to new tasks with minimal adjustments. Initially developed for NLP and widely used in LLMs like GPT-3 Brown et al. (2020), it has since extended to vision-language models. Techniques like prompt tuning Lester et al. (2021) and prefix tuning show that large models can be efficiently adapted using lightweight prompt modifications without full retraining.
In Vision Language Models like CLIP, prompt learning has been particularly successful in enhancing the zero-shot and few-shot learning capabilities of the model Bang et al. (2024); Li et al. (2024b). Language prompts enable the model to better align textual and visual representations, improving performance on tasks like cross-modal retrieval. KgCoOp Yao et al. (2023) enhances language prompts by reducing the discrepancy between learned and handcrafted prompts, thereby improving performance in unseen classes. Vision prompts, on the other hand, modify the image encoding process, allowing the model to adapt to visual-specific tasks without fine-tuning the entire networkChen et al. (2023). Besides, MaPLe khattak et al. (2023a) introduces multi-modal prompt learning, combining language and vision prompt learning. These methods have shown promising results when applied to vision-language models, improving the model’s performance on tasks such as zero-shot image classification and visual question answering. However, the methods above did not provide a comprehensive study of prompt learning in FL. This paper explores both language and vision prompt learning of VLM in FL, where data heterogeneity shows a strong influence on model performance.
2.3 Federated Prompt Learning
Recent research has explored integrating prompt learning into federated systems to collaboratively train shared prompts, enhancing the generalization ability for specific tasks Lu et al. (2023); Guo et al. (2023b); Su et al. (2024); Wei et al. (2023); Gu et al. (2023). For instance, FedPR Feng et al. (2023) leverages federated visual prompts within the null space, while Li et al. (2023) highlights the role of visual prompts in balancing the trade-off between privacy and utility while minimizing privacy budget consumption. In addition to visual prompts, several studies have investigated textual modality adaptation. PromptFL Guo et al. (2023b) focuses on learning a shared text prompt, whereas FedTPG Qiu et al. (2024) introduces a unified prompt generation framework to coordinate prompt learning across clients. Moreover, personalized prompt learning has gained attention, with recent works exploring client-specific prompt optimization strategies to address data heterogeneity Yang et al. (2023); Guo et al. (2023a); Su et al. (2024). FedOTP Li et al. (2024a) proposes an efficient collaborative prompt learning approach that enables each client to capture distinct category characteristics individually. Nevertheless, these methods do not comprehensively test the performance of different prompt learning under various federated settings and prompt configurations.
3 Method Details
We briefly review the foundations of CLIP and introduce the details of prompt learning of CLIP in federated settings.
3.1 Foundations of CLIP
Contrastive Language-Image Pretraining (CLIP) revolutionized multi-modal representation learning by aligning visual and textual data through contrastive learning. Given a dataset of paired image-text examples $D=\left\{(I_{k},T_{k})\right\}_{k=1}^{|D|}$, the goal of CLIP is to encourage semantic alignment between image and text embeddings while repelling unpaired samples. The architecture of CLIP includes a visual encoder $f_{v}$ and a text encoder $f_{t}$,which transform an image $I_{k}$ and a corresponding text $T_{k}$ into normalized feature embeddings $\mathbf{z}_{k}=f_{v}(I_{k})$ and $\mathbf{w}_{k}=f_{t}(T_{k})$ , respectively.
The training objective employs a symmetric contrastive loss based on the InfoNCE Oord et al. (2018) framework. For an image embedding $\mathbf{z}_{k}=f_{v}(I_{k})$, the image-to-text contrastive loss is defined as:
$\mathcal{L}_{I\to T}(\mathbf{z},\mathbf{w})=-\log\frac{\exp(\mathbf{z}_{k}\cdot\mathbf{w}_{k}/\tau)}{\sum_{b=1}^{|B|}\exp(\mathbf{z}_{k}\cdot\mathbf{w}_{b}/\tau)},$ (1)
where $\tau$ is a learnable temperature parameter controlling distribution scaling, $\cdot$ denotes the dot product, and $B$ represents the batch size.
Similarly, the text-to-image contrastive loss, anchored on a text embedding $\mathbf{z}_{k}$, is formulated as:
$\mathcal{L}_{T\to I}(\mathbf{z},\mathbf{w})=-\log\frac{\exp(\mathbf{w}_{k}\cdot\mathbf{z}_{k}/\tau)}{\sum_{b=1}^{|B|}\exp(\mathbf{w}_{k}\cdot\mathbf{z}_{b}/\tau)}.$ (2)
The overall CLIP loss combines these two components symmetrically:
$\mathcal{L}(\mathbf{z},\mathbf{w})=\frac{1}{2}(\mathcal{L}_{I\to T}+\mathcal{L}_{T\to I})$ (3)
3.2 Prompt learning of CLIP in Federated Learning
In this section, we describe how both language and vision prompt learning techniques are adapted for CLIP in FL.
3.2.1 Language Prompt
In Language Prompt Learning (LPT), we utilize a learnable vector $\mathbf{P}_{t}$, which is combined with the class token for the text input, following the approach proposed in Zhou et al. (2022). Specifically, the text input $T_{k}$ is augmented with a series of prompt vectors $\{v_{1},v_{2},\cdots,v_{L}\}$ that are learned during training, where $L$ is the number of tokens in the prompt. The augmented text input to the CLIP text encoder is:
$\tilde{T}_{k}=[v_{1},v_{2},\cdots,v_{L},\texttt{CLASS}],$ (4)
where CLASS is a fixed token that represents the class information (or task-specific information). The input $\tilde{T}_{k}$ is then passed through the text encoder $f_{t}$, resulting in the modified text embedding $\tilde{\mathbf{w}}_{t}$:
$\tilde{\mathbf{w}}_{k}=f_{t}(\tilde{T}_{k}).$ (5)
The $\tilde{\mathbf{w}}_{k}$ is normalized before being used in downstream tasks. The use of learnable text prompts allows the model to adapt to domain-specific language and task-specific variations, improving performance in tasks where text data is heterogeneous across clients.
3.2.2 Vision Prompt
For Vision Prompt Learning (VPT), we follow the method in Jia et al. (2022), where a series of learnable vectors $\{u_{1},u_{2},\cdots,u_{L}\}$ are inserted between the class token CLS and the image patch embeddings $\mathbf{E}$ for the image encoder. The vision prompt $\mathbf{P}_{v}$ is thus represented as a sequence of these learnable vectors, which are concatenated with the image embeddings. The input to the image encoder is:
$\tilde{I}_{k}=[\mathtt{CLS},u_{1},u_{2}\cdots,u_{L},\mathbf{E}].$ (6)
The modified image $\tilde{I}_{k}$ is passed through the visual encoder $f_{v}$, yielding the image embedding $\tilde{\mathbf{z}}_{k}$:
$\tilde{\mathbf{z}}_{k}=f_{v}(\tilde{I}_{k}).$ (7)
Similar to the text embeddings, the resulting image embedding $\tilde{\mathbf{z}}_{k}$ is normalized before being used in downstream tasks. This modification helps the model focus on relevant features and adapt to the visual characteristics specific to each client’s data distribution.
3.2.3 Federated Prompt Learning
In Federated Prompt Learning, the clients collaboratively learn shared prompt modules while maintaining the privacy of their local data. Following typical Federated Learning setup McMahan et al. (2017); Li et al. (2020c, 2021a); Huang et al. (2023b) , we assume that there are $M$ clients (indexed by $M$). Each client $m$ holds its own dataset ${D}^{m}$ and optimizes its local prompts $\mathbf{P}^{m}$ ($\mathbf{P}_{t}$ or $\mathbf{P}_{v}$). The goal is to aggregate the local prompts across clients into a global prompt $\mathbf{P}^{g}$, which is then used to update the global model. The process can be divided into three steps: distribution, optimization, and aggregation.
[…]
where $N^{m}$ is the number of samples at client $m$, and $N=\sum_{m}^{M}N^{m}$ is the total number of samples across all clients. The global prompt $\mathbf{P}^{g}$ is then sent back to all clients, and the process is repeated iteratively. Through this federated learning setup, clients collaboratively adapt the shared prompts to their local data distributions, while the global model maintains privacy and benefits from the collective knowledge.
4 Experiments Setup
4.1 Datasets
Following Li et al. (2021a); Huang et al. (2023b), we comprehensively evaluate the federated prompt learning with CLIP on the following four datasets.
  • Cifar-100 Krizhevsky and Hinton (2009) is a famous classification dataset, containing 32 $\times$ 32 images of 100 classes. Training and validating sets are composed of $50,\!000$ and $10,\!000$ images.
  • DomainNet Peng et al. (2019) contains $6$ domains: Sketch (S), real (R), painting (P), infograph (I), quickdraw (Q), and clipart (C) with $365$ classes. It consists $48k-172k$ images per domain.
  • Office-Home Venkateswara et al. (2017) is a large classification dataset with $65$ classes and includes four different domains: Art (Ar), Clipart (C), Product (P), and Real World (RW).
  • Office31 Saenko et al. (2010) has 31 classification number in three domains: Amazon (Am), DSLR (D), and Webcam (W). It consists of common objects in office scenarios, such as laptops, keyboards, and, file cabinets.
4.2 Implementation Details
We present the experimental setup from three key aspects as follows:
  • Network Structure: We employ the pre-trained CLIP model Radford et al. (2021) with a ViT-B/16 image encoder backbone, using the official implementation from the repository [^1] https://github.com/openai/CLIP. For prompt construction, both for the text and visual encoders, we refer to the open-source implementations of CoOPZhou et al. (2022), Jia et al. (2022), and PromptSRC Khattak et al. (2023b) available at [^2] https://github.com/KaiyangZhou/CoOp, [^3] https://github.com/KMnP/vpt, [^4] https://github.com/muzairkhattak/PromptSRC. The prompt length for both visual and textual prompts is set to $L=16$. Visual prompts $\mathbf{P}_{v}$ are inserted into the first 9 layers of the transformer. Both textual and visual prompts are randomly initialized with the normal distribution.
  • Client Scale Construction: For datasets Office31 Saenko et al. (2010), Office-Home Venkateswara et al. (2017), and DomainNet Peng et al. (2019), which include multiple domains, we design the number of participants to be twice the number of domains. For example, we set the participant size to $K=6$ for Office31. For Cifar-100, we set the client scale to $K=10$.
  • Training Settings: To ensure a fair comparison, we follow Huang et al. (2023b) and set the communication epoch $T=50$ and the local update round $E=1$ for all settings. The training batch size is $32$, and we use SGD as the local update optimizer. The corresponding weight decay is $\eta=1e-5$ and momentum is set to $0.9$. The local client learning rate is $0.001$ for all scenarios. To mimic the label skew, we follow the common experiments setting and use Dirichlet sampling Balakrishnan et al. (2019). We fix the random seed at 00 to ensure reproducibility and run the experiments on NVIDIA 3090.
5 Results and Analysis
Based on the above settings, we conduct extensive experiments to explore different types of prompt learning with CLIP under federated learning settings in detail.
  Cifar100
Methods $\beta=0.3$ $\beta=0.5$ $\beta=1.0$ $\beta=5.0$
ZS-CLIP 64.88
VPT 78.74↑13.86 79.44↑14.56 79.38↑14.50 79.76↑14.88
LPT 75.24↑10.36 75.13↑10.25 76.05↑11.17 76.53↑11.65
Table 1: Accuracy of different prompts learning in Label Skew scenarios. Please refer to Sec. 5.1.
  Office-Home     Office31     Domainnet
Methods A C PR RW     AVG     AM D W     AVG     C I P Q R S     AVG
ZS-CLIP 84.30 66.17 89.18 89.66     82.32     80.96 72.45 74.05     75.82     87.69 69.66 79.77 28.11 91.91 84.82     73.66
VPT 84.13 75.23 93.84 92.14     86.34     87.26 90.00 92.53     89.93     88.93 75.34 82.57 48.83 93.45 87.44     79.43
[…]
Table 2: Accuracy of different prompts learning in Domain Shift scenarios. See details in Sec. 5.1
(a) Mean Value of Cifar-100
(b) Single Client of Cifar-100
(c) Mean Value of Office-Home
(d) Single Client of Office-Home
Figure 2: Similarities of Optimization Directions between clients and global server in Cifar-100 and Office-Home in the training. Fig. 2(a) and Fig. 2(c) illustrate the mean value of the similarity of optimization direction between all clients and the global server, while the shadows indicate the standard deviation. Fig. 2(b) and Fig. 2(d) are the selected clients. Please refer to Sec. 5.1.
Figure 3: Effects of different Communication Epochs. Please refer to Sec. 5.2
(a) VPT in Cifar-100
(b) LPT in Cifar-100
(c) VPT in Office-Home
(d) LPT in Office-Home
Figure 4: Similarities of Optimization Directions among clients in Cifar-100 and Office-Home in the training epoch. Fig. 4(a) and Fig. 4(b) are the similarity between clients of VPT and LPT in Cifar-100 respectively, while Fig. 4(c) and Fig. 4(d) stand for Office-Home similarly. See details in Sec. 5.1.
5.1 Behavior Discrepancy between LPT and VPT in Label Skew and Domain Shift
To fully explore Language Prompt Learning (LPT) and Vision Prompt Learning (VPT) with CLIP in federated learning settings, we conduct both experiments in the Label Skew and Domain Shift scenarios.
Label Shift Scenarios. Applying Dirichlet sampling, we evaluate VPT and LPT in Cifar-100 with different degrees of Label Skew and the results are shown in Tab. 1. VPT outperforms LPT obviously in all settings.
Domain Shift Scenarios. We conduct experiments on Office-Home, Office31 and DomainNet as for Domain Shift scenarios. The results in Tab. 2 indicate that LPT works better than VPT in most domains.
Since the results in the two scenarios are opposite, the key problem arises: What makes the differences of results in Label Skew and Domain Shift?
To address this problem, we further step in the training process in both scenarios. In concrete, we calculate the similarities of optimization directions of prompt learning between clients and the global server during training, and the visualization results are shown in Fig. 2. We can observe that in both Label Skew and Domain Shift scenarios, clients tend to optimize more similarly with the global server during the training of VPT. Besides, we also record the similarities among clients in Cifar-100 and Office-Home. We select the results of a single communication epoch and visualize them in Fig. 4. Likewise, the optimization directions among clients are more similar in VPT than in LPT.
Our experiments reveal that under label skew conditions, VPT exhibits significantly higher consistency in optimization directions among clients and between clients and the global server. This consistency helps align the visual features across disparate client data, effectively mapping the textual cues into a unified visual domain. In contrast, LPT shows more variability in its optimization paths, leading to inconsistent mapping and thus poorer performance when label distributions are imbalanced. Under domain shift scenarios, the optimization directions in LPT are less similar across clients, which appears to allow each client to better adapt to its own domain-specific data. This individualized adaptation enables LPT to map the textual features into diverse visual domains more effectively, yielding superior performance. Conversely, the high consistency of VPT in optimization directions tends to ignore the unique characteristics of each domain, thereby limiting its adaptability and resulting in lower performance in heterogeneous domain settings.
Regarding Q1 in Sec. 1, we can obtain the answer: A1: There exists Behavior Discrepancy between LPT and VPT under label skew and domain shift. We restate this key phenomenon as follows: {mdframed}[backgroundcolor=cadetblue!10, linewidth=0.8pt, linecolor=cadetblue!80, roundcorner=5pt] Label Skew Prompt Consistency:
For label skew, VPT exhibits consistent optimization across clients and with the global server, leading to unified visual mapping.
{mdframed}
[backgroundcolor=cadetblue!10, linewidth=0.8pt, linecolor=cadetblue!80, roundcorner=5pt] Domain Shift Prompt Discrepancy:
For domain shift, diverse optimization in LPT enables better adaptation to distinct domains.
5.2 Robustness of Federated Prompt Learning Under Various Settings
(a) Cifar-100
(b) Office-Home
Figure 5: Effects of different Aggregation Ways. E is short for Equal, while W is short for Weighted. The dotted line represents the effect of CLIP with zero-shot inference. See details in Sec. 5.2.
To comprehensively evaluate the robustness of FPL, we conduct extensive experiments under different federated learning settings and prompt configurations, analyzing their impact on model performance. The following subsections detail our findings across communication epochs, aggregation strategies, client scales, and prompt lengths.
Communication Epochs. We assess the effect of varying the number of communication epochs on performance by conducting experiments on Cifar-100 and Office-Home. As shown in Fig. 3, increasing the number of communication rounds leads to only marginal improvements in accuracy, while still maintaining a performance advantage over direct zero-shot inference of CLIP. This suggests that Federated Prompt Learning is relatively robust to the number of training epochs, indicating that beyond a certain point, additional training contributes little to accuracy gains.
From a practical standpoint, this observation highlights an important trade-off: while increasing training rounds may slightly refine the model, it also incurs additional computational and communication costs. Therefore, in real-world applications, it is crucial to determine an optimal balance between training cost and performance improvement.
  Office-Home     Office31     Domainnet
Methods A C PR RW     AVG     AM D W     AVG     C I P Q R S     AVG
ZS-CLIP 84.30 66.17 89.18 89.66     82.32     80.96 72.45 74.05     75.82     87.69 69.66 79.77 28.11 91.91 84.82     73.66
Lable skew: $\beta=0.5$
[…]
LPT 86.45 75.99 93.94 93.17     87.39     88.15 93.47 94.43     92.02     88.89 73.75 84.89 43.74 93.64 87.00     78.65
[…]
Lable skew: $\beta=1.0$
[…]
LPT 87.44 74.01 94.66 92.53     87.16     87.97 91.63 95.44     91.68     89.74 75.32 83.41 45.94 93.76 87.03     79.20
[…]
Table 3: Performance of different prompt learning in complex scenarios containing both Label Skew and Domain Shift. Please refer to Sec. 5.3
Aggregation Ways. To investigate the impact of aggregation strategies in FL settings, we compare two common methods: Weighted and Equal. While Weighted aggregates client models by assigning weights based on the size of each client’s dataset, Equal assumes the weights of clients are equal.
Our experiments on Cifar-100 and Office-Home shown in Fig. 5 reveal an interesting pattern: In label skew scenarios, weighted aggregation leads to better performance, as larger clients contribute more meaningfully to the global prompt optimization. In domain shift scenarios, equal aggregation performs better, as it prevents overfitting to the data-rich clients and helps maintain diversity across domains. However, it is important to note that the overall difference in performance between these aggregation methods is relatively small. More importantly, both methods still outperform CLIP’s zero-shot inference, reinforcing the robustness of federated prompt learning to different aggregation strategies. ggests that while aggregation choice may slightly influence results, FPL remains stable and effective across different settings.
(a) Cifar-100
(b) Office-Home
Figure 6: Effects of different Scale of Clients. The dotted line represents the effect of CLIP with zero-shot inference. Please refer to Sec. 5.2.
Scale of Clients. To evaluate the impact of client scale in federated learning settings, we conduct experiments on Cifar-100 under label skew conditions and DomainNet under domain shift conditions. The results are presented in Fig. 6.
Our findings indicate that in label skew scenarios, model accuracy gradually decreases as the number of clients increases. This is likely due to the increasing divergence in class distributions across clients, making it more challenging to learn a globally consistent prompt representation. In contrast, under domain shift scenarios, there is no clear pattern in accuracy fluctuations as the client count changes. Nevertheless, across both settings, the variation in client scale does not significantly degrade model performance, and in all cases, FPL consistently outperforms CLIP’s zero-shot inference. This demonstrates the robustness of FPL to different client scales, making it a reliable approach even in privacy-friendly environments with varying numbers of participants.
  Cifar100
Length $\beta=0.3$ $\beta=0.5$ $\beta=1.0$ $\beta=5.0$
ZS-CLIP 64.88
[…]
32 79.95↑15.07 79.87↑14.99 80.49↑15.61 80.32↑15.44
64 80.01↑15.13 80.10↑15.22 80.58↑15.70 80.92↑16.04
128 80.32↑15.44 80.37↑15.49 81.21↑16.33 81.56↑16.68
  Office-Home
Length $\beta=0.3$ $\beta=0.5$ $\beta=1.0$ $\beta=5.0$
ZS-CLIP 82.32
[…]
32 86.09↑3.77 86.38↑4.06 86.47↑4.15 86.48↑4.16
64 86.47↑4.15 86.52↑4.20 86.36↑4.04 86.54↑4.22
128 86.62↑4.30 86.74↑4.42 86.22↑3.90 86.85↑4.53
Table 4: Effects of different Length of Prompt on VPT in Cifar-100 and Office-Home. See details in Sec. 5.2.
Prompt Lengths. We further examine the effect of prompt length as a critical parameter in prompt-based learning. We conduct experiments on Cifar-100 and Office-Home, using visual prompt tuning (VPT), and present the results in Tab. 4. Key observations include: (a) Across multiple label skew settings, increasing the prompt length (e.g., doubling its size) generally leads to a slight improvement in accuracy. (b) Even with minimal prompt tuning, FPL consistently surpasses CLIP’s zero-shot inference, demonstrating its robustness to variations in prompt length. While longer prompts can provide slight accuracy gains, they also increase the computational and memory overhead required for training. In real-world deployment, an optimal trade-off must be considered, balancing improved performance with resource constraints.
The above experiments collectively demonstrate that changes in communication epochs, aggregation strategies, scale of clients, and prompt length lead to minor variations in performance. In all cases, FPL outperforms CLIP’s zero-shot inference, reinforcing its adaptability and effectiveness in heterogeneous federated learning environments. We can answer Q2 in Sec. 1 that: A2: Federated Prompt Learning remains robust under different settings.
5.3 Collaboration Effect in Complex Scenarios
In real-world environments, label skew and domain shift often coexist, creating highly complex learning conditions for federated prompt learning. To address this, we explore strategies to enhance Federated Prompt Learning (FPL) under such challenging conditions. Specifically, we investigate whether combining both vision prompt learning (VPT) and language prompt learning (LPT) can improve performance when computational resources allow. This combined approach, referred to as VLPT (Vision-Language Prompt Tuning), jointly leverages both textual and visual prompts to better align multimodal representations.
We evaluate VLPT across multiple datasets, including Office-Home, Office31, and DomainNet, where we introduce label skew on top of the existing domain shift settings. The experimental results are presented in Tab. 3 and several key findings are recapped below:
  • In all tested complex scenarios, VLPT consistently achieves the best performance, outperforming both standalone VPT and LPT across datasets.
  • The results in Sec. 5.1 suggest that LPT is more effective under domain shift, whereas VPT performs better in label skew scenarios. By combining both, VLPT benefits from their complementary strengths, improving the overall generalization capability.
  • As seen in Tab. 3, VLPT provides noticeable improvements across diverse domains, particularly in DomainNet, where domain shifts are more severe.
Practical Implications. As for Q3 in Sec. 1, we have the conclusion that: A3: In real-world federated learning applications, where both label skew and domain shift are frequently common, VLPT provides an effective solution for boosting model performance. If computational resources permit, utilizing both vision and language prompts simultaneously can significantly enhance the adaptability and overall robustness in federated learning settings.
6 Conclusion and Future Work
In this paper, we present a comprehensive empirical study on prompt learning for Vision Language Models (VLMs) in federated settings. Based on the experiments above, we can draw the key conclusions corresponding to the main research objectives as below:
  • [Uncaptioned image]
    A1: Under label skew and domain shift, LPT and VPT show significant Behavior Discrepancies. LPT excels in domain shift scenarios, while VPT works better in label skew since more similar updating directions are not always the “antidote” to data heterogeneity.
  • [Uncaptioned image]
    A2: Federated prompt learning exhibits Robustness Patterns when federated and prompt settings vary, such as communication epochs, aggregation ways, client scales and prompt lengths.
  • [Uncaptioned image]
    A3: Integrating LPT and VPT helps to adapt in both branches of language and vision, thus showing Collaboration Effect in complex scenarios containing label skew and domain shift.
While our findings highlight the feasibility and advantages of prompt learning in federated environments, there remain several avenues for future exploration. Current limitations in training time and computational resources restrict the scalability of our experiments, and further research is needed to enhance the efficiency, personalization, and adaptability of federated prompt learning.
Contribution Statement
Equal contribution. Corresponding author.
Acknowledgments
This work is supported by the National Key Research and Development Program of China (2024YFC3308400), National Natural Science Foundation of China under Grant (62032016, 62361166629, 62176188, 623B2080), Zhongguancun Laboratory, and the Wuhan University Undergraduate Innovation Research Fund Project.

### v2

@@
 The process can be divided into three steps: distribution, optimization, and aggregation.
-[…]
-[…]
-[…]
+$\mathbf{P}^{m}\leftarrow{}\mathbf{P}^{g}$ $\text{Distribution},$ (8)
+$\text{arg}\min_{\mathbf{P}^{m}}\mathbb{E}_{(x,{y})\sim D^{m}}\,\mathcal{L}(\mathbf{{z}},\mathbf{\tilde{w}})/\mathcal{L}(\mathbf{\tilde{z}},\mathbf{{w}})$ $\text{Optimization},$
+$\mathbf{P}^{g}=\sum_{m}^{M}\frac{N^{m}}{N}\mathbf{P}^{m}$ $\text{Aggregation},$
 where $N^{m}$ is the number of samples at client $m$, and $N=\sum_{m}^{M}N^{m}$ is the total number of samples across all clients.
@@
   • Cifar-100 Krizhevsky and Hinton (2009) is a famous classification dataset, containing 32 $\times$ 32 images of 100 classes.
-Training and validating sets are composed of $50,\!000$ and $10,\!000$ images.
-  • DomainNet Peng et al. (2019) contains $6$ domains: Sketch (S), real (R), painting (P), infograph (I), quickdraw (Q), and clipart (C) with $365$ classes. It consists $48k-172k$ images per domain.
+Training and validating sets are composed of 50,00050,\!000 and 10,00010,\!000 images.
+  • DomainNet Peng et al.
+(2019) contains $6$ domains: Sketch (S), real (R), painting (P), infograph (I), quickdraw (Q), and clipart (C) with $365$ classes.
+It consists 48k172k48k-172k images per domain.
   • Office-Home Venkateswara et al. (2017) is a large classification dataset with $65$ classes and includes four different domains: Art (Ar), Clipart (C), Product (P), and Real World (RW).
@@
 For prompt construction, both for the text and visual encoders, we refer to the open-source implementations of CoOPZhou et al.
-(2022), Jia et al.
+(2022), VPTJia et al.
 (2022), and PromptSRC Khattak et al.
@@
 (2019).
-We fix the random seed at 00 to ensure reproducibility and run the experiments on NVIDIA 3090.
+We fix the random seed at 0 to ensure reproducibility and run the experiments on NVIDIA 3090.
 5 Results and Analysis
@@
 1, we can obtain the answer: A1: There exists Behavior Discrepancy between LPT and VPT under label skew and domain shift.
-We restate this key phenomenon as follows: {mdframed}[backgroundcolor=cadetblue!10, linewidth=0.8pt, linecolor=cadetblue!80, roundcorner=5pt] Label Skew Prompt Consistency:
-For label skew, VPT exhibits consistent optimization across clients and with the global server, leading to unified visual mapping.
-{mdframed}
-[backgroundcolor=cadetblue!10, linewidth=0.8pt, linecolor=cadetblue!80, roundcorner=5pt] Domain Shift Prompt Discrepancy:
-For domain shift, diverse optimization in LPT enables better adaptation to distinct domains.
+We restate this key phenomenon as follows: Label Skew Prompt Consistency: For label skew, VPT exhibits consistent optimization across clients and with the global server, leading to unified visual mapping.
+Domain Shift Prompt Discrepancy: For domain shift, diverse optimization in LPT enables better adaptation to distinct domains.
 5.2 Robustness of Federated Prompt Learning Under Various Settings
@@
 Lable skew: $\beta=0.5$
-[…]
+VPT 85.08 74.89 93.21 92.25     86.62     85.66 90.82 89.62     88.70     88.86 74.59 83.41 49.70 93.25 86.74     79.43
 LPT 86.45 75.99 93.94 93.17     87.39     88.15 93.47 94.43     92.02     88.89 73.75 84.89 43.74 93.64 87.00     78.65
@@
 Lable skew: $\beta=1.0$
-[…]
+VPT 84.92 73.78 93.78 91.93     86.10     85.98 90.20 89.49     88.56     88.95 75.27 82.51 49.80 93.35 86.61     79.41
 LPT 87.44 74.01 94.66 92.53     87.16     87.97 91.63 95.44     91.68     89.74 75.32 83.41 45.94 93.76 87.03     79.20
@@
 Our experiments on Cifar-100 and Office-Home shown in Fig.
-5 reveal an interesting pattern: In label skew scenarios, weighted aggregation leads to better performance, as larger clients contribute more meaningfully to the global prompt optimization.
-In domain shift scenarios, equal aggregation performs better, as it prevents overfitting to the data-rich clients and helps maintain diversity across domains.
+5 reveal an interesting pattern: In label skew scenarios, equal aggregation leads to better performance, as treating clients uniformly prevents bias toward label-rich participants and ensures balanced optimization across classes.
+In domain shift scenarios, weighted aggregation performs better, as emphasizing larger domains helps the global model capture the dominant distribution while still retaining cross-domain knowledge.
 However, it is important to note that the overall difference in performance between these aggregation methods is relatively small.
@@
 ZS-CLIP 64.88
-[…]
+16 78.74↑13.86 79.44↑14.56 79.38↑14.50 79.76↑14.88
 32 79.95↑15.07 79.87↑14.99 80.49↑15.61 80.32↑15.44
@@
 ZS-CLIP 82.32
-[…]
+16 86.11↑3.79 86.36↑4.04 86.10↑3.78 86.71↑4.39
 32 86.09↑3.77 86.38↑4.06 86.47↑4.15 86.48↑4.16